# Lan Wang1,2∗        Yujia Chen2        Du Tran2       Vishnu Naresh Boddeti1       Wen-Sheng Chu2

### v2

@@ Abstract
Long video understanding presents challenges due to the inherent high computational complexity and redundant temporal information. An effective representation for long videos must process such redundancy efficiently while preserving essential contents for downstream tasks. This paper introduces SEmantic Attention Learning (SEAL), a novel unified representation for long videos. To reduce computational complexity, long videos are decomposed into three distinct types of semantic entities: scenes, objects, and actions, allowing models to operate on a handful of entities rather than a large number of frames or pixels. To further address redundancy, we propose an attention learning module that balances token relevance with diversity, formulated as a subset selection optimization problem. Our representation is versatile, enabling applications across various long video understanding tasks. Extensive experiments show that SEAL significantly outperforms state-of-the-art methods in video question answering and temporal grounding tasks and benchmarks including LVBench, MovieChat-1K, and Ego4D.
@@ 1 Introduction
This work was done as Lan Wang’s internship project at Google.
@@
State-of-the-art video understanding models perform well on short video tasks such as video classification [14, 8] and action detection [13, 11], which involve videos lasting from a few seconds to a few minutes, but may fall short on hour-long videos [40, 41]. In contrast, a 10-year-old child can watch a full-length movie (one to two hours) and answer questions about it at various levels. This human-versus-AI perception gap emphasizes the unique challenges in long video understanding for AI models, which include (1) Increased complexity: Long videos require more computation and memory than current hardware can support for training or inference, (2) Temporal redundancy: Scenes and objects often change slowly, leading to significant redundancy over time, and (3) Cross-task generalization: A robust representation must be applicable to various tasks, from detailed fact-finding to high-level reasoning. These challenges are difficult for AI models, yet seem natural for a 10-year-old child.
@@
How do human brains process long videos, particularly to address the above-mentioned challenges? First, it is evident that the brain selectively attends to new information rather than processing every pixel or frame [4, 23], thus effectively managing temporal redundancy. Second, humans process videos in an online fashion, continuously updating their understanding and memories as they watch, rather than waiting until the end to process the entire video. This “online” knowledge, along with the selective attention, allows the brain to handle the complexity of long videos efficiently. Finally, the attention may shift depending on whether we know the questions to be asked afterwards. Without specific questions, a child may naturally focus on memorable or exciting moments. However, if questions are known in advance, attention may shift to seek relevant details while still retaining a broad understanding. This suggests that effective representations should balance between task-specific focus and holistic understanding of the video to enable cross-task generalization.
@@
Inspired by how humans process long videos, we introduce SEmantic Attention Learning (SEAL), a novel unified representation designed to tackle the three key challenges in long video understanding. SEAL consists of two main steps: Semantic Decomposition and Attention Learning. In Semantic Decomposition, long videos are first decomposed into semantic entities such as scenes, objects, and actions which are then treated as tokens. While the scene and object tokens represent “static” content of the video and assumed not to change rapidly, the action tokens are designed to capture the “dynamic” and fast-changing moments of the video. We note that these semantic tokens are sufficient enough for answering to “where”, “what”, or “how” questions about the videos. This semantic decomposition helps to efficiently reduce the complexity as AI models can now operate on a much smaller number of tokens instead of a large number of pixels or frames. Figure 1(a) illustrates a conventional uniformly sampled video $\mathcal{V}$, where redundant frames create cluttered visual information that hinders effective analysis for both AI models and humans. Figure 1(b) shows our semantic decomposition, breaking down the video into distinct scenes, “static” objects, and “dynamic” actions (or events). In the second step, we introduce Attention Learning, formulated as a subset selection problem which maximizes the query-relevance and maintains a good token-diversity. We note that our attention learning module not only address the redundancy by enforcing token diversity, but also improve cross-task generalization via maximizing query-relevance. Figure 1(c) demonstrates our attention learning module with four different queries visualized with the selected tokens capturing their relevance to queries and diverse video content. Finally, SEAL is designed to work for both global and streaming modes, enabling it to process arbitrarily long videos. Extensive ablations and experiments confirm SEAL’s superior performance over existing methods. Our key contributions include:
@@
  • We introduce SEAL, a novel unified representation for long videos by decomposing them into semantic tokens, namely scenes, objects, and actions.
@@
  • SEAL significantly outperforms state-of-the-art methods on various long video understanding tasks and benchmarks including: video QA (MovieChat-1K [34], LVBench [40]), and egocentric video grounding (Ego4D [9]).
@@ 3.1 Semantic Decomposition of Long Videos
The main challenge in long video representation lies in capturing diverse content within limited memory. Conventional methods resort to frame sampling [44, 42] or maintaining a memory bank that merges similar frames [34, 35]. However, these approaches can vary greatly across tasks. We propose a novel decomposition approach that structures long videos into three distinct token types representing different type of semantic entities: (1) Scene tokens $\mathbf{T}_{\text{scene}}$ capture background context, providing essential cues about the environment. (2) Action tokens $\mathbf{T}_{\text{action}}$ represent moving elements, focusing on temporal information such as motions, activities, or events. (3) Object tokens $\mathbf{T}_{\text{object}}$, highlight key static elements relevant to specific tasks. With this structured tokenization, we create a unified, compact, task-agnostic representation that minimizes the need for redundant and dense video storage while preserving a comprehensive understanding of long video content.
@@
Action Tokens. The purpose of the action token is to capture temporal information of moving objects such as low-level motions, activities, and events. We begin by using a class-agnostic object tracker, e.g., SAM-2 [32], to extract multiple initial dynamic tracklets. Tracklets shorter than $L_{\text{min}}$ are discarded, while those longer than $L_{\text{max}}$ are split into multiple trackets with length of $L_{\text{max}}$. For each tracklet, we take the spatial union of bounding boxes across frames, allowing the dynamic token to capture not only motion information but also the spatial movement of people and objects. This process yields $N_{\text{tracklet}}$ tracklets, denoted as $\mathbf{T}_{\text{action}}=\{\tau_{\text{dynamic}}^{i}\}_{i=1}^{N_{\text{tracklet}}}$
Object Tokens. For the object tokens, we utilize a class-agnostic grouping method such as SAM [15] to generate masks for all objects in each frame. This class-agnostic segmentation approach enables comprehensive object information capture. Specifically, we apply this process on $N_{\text{key}}$ key frames. On each of the $N_{\text{key}}$ frames, we apply SAM to obtain in total $N_{\text{object}}$ object masks, which we denote as $\mathbf{T}_{\text{object}}=\{\mathbf{M}_{\text{object}}^{i}\}_{i=1}^{N_{\text{object}}}$. Depending on the specific task, we select different key frames, which will be detailed in the following sections.
@@ 3.2 Attention Learning
$\begin{split}T_{s}^{*}&=\max_{T_{s}\subset T_{G}}F_{s}(T_{s}|T_{G},q)\\ &=\max_{T_{s}\subset T_{G}}\alpha\sum_{t_{s}\in T_{s}}R(t_{s},q)+(1-\alpha)\sum_{t_{i},t_{j}\in T_{s}}\frac{1}{S(t_{i},t_{j})}.\end{split}$
@@ 3.4 Prediction Heads
The regression and classification heads are optimized using an IoU distance and a focal loss as used in [26]. The final moment is calculated as:
$t_{start}^{i},t_{end}^{i}=(t_{i}-d_{\text{start}}^{i})\times L_{V},(t_{i}+d_{\text{end}}^{i})\times L_{V},$ (4)
where $L_{V}$ is the length of the input video. The proposed unified representation enables efficient handling of long video sequences and allows for localization related tasks.
@@ 4.2 Comparison with State-of-the-arts
Table 4: The effects of different types of tokens on LVbench. Using any single type of tokens outperforms the random baseline validates the effectiveness of our semantic tokens. Using all three types of tokens brings the best performance.
@@
4.3 Ablation Study
Semantic decomposition. We analyze the impact of the three proposed token types on LVBench are present results in Table 4. To prevent Large Language Models (LLMs) from leveraging prior knowledge to “guess” answers using only the text query, we present the baseline that uses random token inputs, isolating the LLM’s performance. Results indicate that using each token type individually significantly improves performance, highlighting the value of each token’s contribution. Notably, scene tokens provide the most substantial improvement, aligning with prior research [37, 2, 41] that emphasizes frame-based scene tokens. Furthermore, different token combinations yield additional gains: adding action tokens enhances Key Information Retrieval (KIR), Reasoning (Rea) and Temporal Grounding (TG) tasks by capturing temporal dynamics in the queries, while object tokens boost Entity Recognition (ER) performance by retaining detailed object-specific information. We observe high variance in the Summarization (Sum) accuracy due to the small number of questions in this category and the randomness of LLM. Similar improvements have been observed on the MovieChat-1K in Table 4.2, where the accuracy improved by 10.45% with action and object tokens. Ultimately, using all three token types together achieves the highest performance, demonstrating the complementary strengths of scene, action, and object tokens as a unified representation in long video understanding.
We observe that the latency bottleneck mostly comes from the tracker when processing multiple objects. Table 4.3 presents an analysis of different trackers for extracting action tokens. While YOLO+BoT-SORT operates in a much higher frame rate (14 FPS), it lacks open-set detection capabilities. On the other hand, SAM2 can detect all objects within a scene but operates significantly slower (8FPS). Our findings indicate that the proposed action tokens does not requires capturing all the objects in the scene. In practice, YOLO-X pretrained on COCO, with BoT-SORT, effectively captures most of the essential information required for action token extraction, maintaining a balance between performance and efficiency.
Attention learning. Table 4.3 demonstrates the effectiveness of the query-relevance and the token-diversity terms in Section 3.2. When $\alpha=0$, the optimization focuses solely on the diversity term, while $\alpha=1$ prioritize the relevance term. Optimizing only for diversity enables the method to attend to a variety of content in long videos, achieving reasonable results comparable to other 34B state-of-the-art approaches in Table 1, particularly for tasks requiring extensive contextual information, such as Summarization (Sum) and Reasoning (Rea). On the other hand, optimizing only for relevance effectively reduces redundancy by sampling tokens that closely align with the query, yielding better performance in tasks like Key Information Retrieval (KIR). However, this approach loses global context and results in weaker summarization performance. Combining both terms achieves the best overall results.
@@ 4.3 Ablation Study
Table 6: Ablation studies of different component on LVBench. SEAL defaults using global mode with YOLO. SAM2 yields comparable results, while YOLO being more efficient. Global mode excels in long-context tasks and streaming mode in temporally-intensive tasks. Combining Relevance and Diversity in attention learning achieves the best performance.
5 Conclusion and Future works
We propose SEAL, a novel unified representation for long video understanding that addresses three key challenges: computational complexity, temporal redundancy, and cross-task generalization. SEAL leverages semantic decomposition to break videos into scene, object, and action tokens, reducing redundancy and enabling efficient processing. It incorporates attention learning to balance query relevance and token diversity, enhancing performance across diverse tasks. SEAL achieves state-of-the-art results on benchmarks like MovieChat-1K, LVBench, and Ego4D-NLQ, demonstrating its versatility and effectiveness for long video understanding.
Limitation: The Attention Learning module is bounded by the memory constraint for the QP solver, making it not fully end-to-end trainable. For future work, we plan to integrate Attention Learning in our streaming mode for full end-to-end learning. Developing special prediction heads to solve causal reasoning [17], where MLLM heads showed limitations, is also an interesting area for future exploration.
Appendix A Further Ablations
@@
Figure 4 shows the performance trends across various categories for different values of $\alpha$. $\alpha=0.9$ achieves the best overall trade-off, reaching peak with the highest overall accuracy of 45.9. Conversely, extreme values like $\alpha=0.0$ or $1.0$ lead to declines in several metrics, highlighting that both diversity and relevance are essential. Therefore, $\alpha=0.9$ is the optimal choice for experiments, delivering peak performance and a well-rounded balance across all categories.
Appendix B Result Analysis
We evaluated the distribution of answers generated by different models, following [40], as shown in Figure 6. The Ground-Truth exhibits a fairly balanced distribution among A, B, C, and D, indicating a well-distributed dataset where no single category is disproportionately represented. MovieChat and LWM shows a dominance of category A, with significantly smaller contributions from other categories, suggesting a lack of diversity in predictions. In Gemini 1.5 Pro, the “Others” category is significantly high, indicating that Gemini 1.5 Pro produces a notable number of unrelated outputs. Our method demonstrates a distribution close to Ground-Truth, showing strong generalization and robustness.

### v3

# Lan Wang1,2        Yujia Chen2        Du Tran2       Vishnu Naresh Boddeti1       Wen-Sheng Chu2

@@ Abstract
Long video understanding presents challenges due to the inherent high computational complexity and redundant temporal information. An effective representation for long videos must efficiently process such redundancy while preserving essential contents for downstream tasks. This paper introduces SEmantic Attention Learning (SEAL), a novel unified representation for long videos. To reduce computational complexity, long videos are decomposed into three distinct types of semantic entities: scenes, objects, and actions, allowing models to operate on a compact set of entities rather than a large number of frames or pixels. To further address redundancy, we propose an attention learning module that balances token relevance with diversity, formulated as a subset selection optimization problem. Our representation is versatile and applicable across various long video understanding tasks. Extensive experiments demonstrate that SEAL significantly outperforms state-of-the-art methods in video question answering and temporal grounding tasks across diverse benchmarks, including LVBench, MovieChat-1K, and Ego4D.
@@ 1 Introduction
State-of-the-art video understanding models excell at short video tasks such as video classification [14, 8], temporal grounding [39] and action detection [13, 11], which involve videos lasting from a few seconds to minutes. However, their performance declines on hour-long videos [41, 42]. In contrast, a 10-year-old child can watch a full-length movie (one to two hours) and effortlessly answer questions at various levels of detail. This disparity between humans and machines emphasizes the foundamental challenges in long video understanding for AI models, including: (1) Increased complexity: Long videos require more computation and memory than current hardware can support for training or inference, (2) Temporal redundancy: Slow-changing scenes and objects introduce significant redundancy, and (3) Cross-task generalization: A robust representation must adapt to various tasks, from fine-grained fact retrieval to high-level reasoning. These challenges, which appear trivial for a child, remain challenging for AI models.
@@
How does the human brain process long videos, particularly in addressing the above-mentioned challenges? First, rather than processing every pixel or frame, the brain selectively attends to new information to efficiently manage temporal redundancy [4, 23]. Second, humans process videos in an online fashion, continuously updating their understanding and memories as they watch, rather than deferring reasoning until the end. This continuous knowledge update, combined with selective attention, allows the brain to efficiently handle the complexity of long videos. Finally, attention dynamically shifts based on context. Without specific guidance, a child may focus on naturally engaging or memorable moments. However, when given specific questions in advance, attention becomes goal-oriented to seek relevant details while maintaining a broad understanding. This suggests effective representations should balance between task-specific focus and holistic understanding of the video to enable cross-task generalization.
@@
Inspired by how humans process long videos, we introduce SEmantic Attention Learning (SEAL), a novel unified representation designed to tackle the three key challenges in long video understanding. SEAL consists of two main steps: Semantic Decomposition and Attention Learning. In Semantic Decomposition, long videos are decomposed into three semantic entities such as scenes, objects, and actions, which are then treated as tokens. While the scene and object tokens represent static content assumed not to change rapidly, the action tokens are designed to capture the dynamic, fast-changing moments of the video. We note that these semantic tokens efficiently encode the essential information needed to answer “where”, “what”, or “how” questions about the videos. This decomposition significantly reduces complexity by allowing AI models to operate on a compact set of tokens instead of raw pixels or frames. Figure 1(a) illustrates a conventional uniformly sampled video $\mathcal{V}$, where redundant frames create cluttered visual information that hinders effective analysis for both models and humans. Figure 1(b) shows our semantic decomposition, breaking down the video into distinct scenes, static objects, and dynamic actions. In Attention Learning, we formulate a subset selection problem that maximizes the query relevance while ensuring token diversity. This step not only mitigates redundancy but also enhances cross-task generalization by prioritizing the most informative tokens. Figure 1(c) illustrates our attention learning module with four different queries and their selected tokens that capture relevance and diverse video content. Finally, SEAL is designed to work for both global and streaming modes, enabling it to process arbitrarily long videos. Extensive ablations and experiments confirm SEAL’s superior performance over existing methods. Our key contributions include:
@@
  • We introduce SEAL, a novel unified representation for long videos by decomposing them into three semantic tokens, namely scenes, objects, and actions.
@@
  • SEALoutperforms state-of-the-art methods on various long video understanding tasks and benchmarks including: video QA (MovieChat-1K [34], LVBench [41]), and egocentric video grounding (Ego4D [9]).
@@ 3.1 Semantic Decomposition of Long Videos
The main challenge in long video representation lies in capturing diverse content within limited memory. Conventional methods resort to frame sampling [45, 43] or maintaining a memory bank that merges similar frames [34, 35]. However, these approaches can vary greatly across tasks. We propose a novel decomposition approach that structures long videos into three distinct token types representing different type of semantic entities: (1) Scene tokens $\mathbf{T}_{\text{scene}}$ capture background context, providing essential cues about the environment. (2) Object tokens $\mathbf{T}_{\text{object}}$, highlight key static elements relevant to specific tasks. (3) Action tokens $\mathbf{T}_{\text{action}}$ represent moving elements, focusing on temporal information such as motions, activities, or events. With this structured tokenization, we create a unified, compact, task-agnostic representation that minimizes the need for redundant and dense video storage while preserving a comprehensive understanding of long video content.
@@
Action Tokens. The purpose of the action token is to capture temporal information of moving objects such as low-level motions, activities, and events. We begin by using a class-agnostic object tracker, e.g., SAM-2 [32], to extract multiple initial dynamic tracklets. Tracklets shorter than $L_{\text{min}}$ are discarded, while those longer than $L_{\text{max}}$ are split into multiple tracklets with length of $L_{\text{max}}$. For each tracklet, we take the spatial union of bounding boxes across frames, allowing the dynamic token to capture not only motion information but also the spatial movement of people and objects. This process yields $N_{\text{tracklet}}$ tracklets, denoted as $\mathbf{T}_{\text{action}}=\{\tau_{\text{dynamic}}^{i}\}_{i=1}^{N_{\text{tracklet}}}$
Object Tokens. For the object tokens, we utilize a class-agnostic grouping method such as SAM [15] to generate masks for all objects in each frame. This class-agnostic segmentation approach enables comprehensive object information capture. Specifically, we apply this process on $N_{\text{key}}$ key frames. On each of the $N_{\text{key}}$ frames, we apply SAM to obtain in total $N_{\text{object}}$ object masks, which we denote as $\mathbf{T}_{\text{object}}=\{\mathbf{M}_{\text{object}}^{i}\}_{i=1}^{N_{\text{object}}}$. Key frame selection varies by task and is detailed in the following sections.
@@ 3.2 Attention Learning
$\begin{split}T_{s}^{*}&=\operatorname*{arg\,max}_{T_{s}\subset T_{G}}F_{s}(T_{s}|T_{G},q)\\ &=\operatorname*{arg\,max}_{T_{s}\subset T_{G}}\alpha\!\!\sum_{t_{s}\in T_{s}}\!\!R(t_{s},q)\!+\!(1\!-\!\alpha)\!\!\!\!\!\sum_{t_{i},t_{j}\in T_{s},i\neq j}\!\frac{1}{S(t_{i},t_{j})}.\end{split}$
@@ 3.4 Prediction Heads
The regression and classification heads are optimized using an IoU distance and a focal loss as used in [26]. The final moment is calculated as $t_{start}^{i},t_{end}^{i}=(t_{i}-d_{\text{start}}^{i})\times L_{V},(t_{i}+d_{\text{end}}^{i})\times L_{V}$, where $L_{V}$ is the length of the input video. The proposed unified representation enables efficient handling of long video sequences and allows for localization related tasks.
@@ 4.2 Comparison with State-of-the-arts
4.3 Ablation Study
Semantic decomposition. We analyze the impact of the three proposed token types on LVBench (Table 4). To prevent LLMs from leveraging prior knowledge to “guess” answers using only the text query, we present the baseline that uses random token inputs, isolating the LLM’s performance. Results indicate that using each token type individually significantly improves performance, highlighting the value of each token’s contribution. Notably, scene tokens provide the most substantial improvement, aligning with prior research [37, 2, 42] that emphasizes frame-based scene tokens. Furthermore, different token combinations yield additional gains: adding action tokens enhances Key Information Retrieval (KIR), Reasoning (Rea) and Temporal Grounding (TG) tasks by capturing temporal dynamics in the queries, while object tokens boost Entity Recognition (ER) performance by retaining detailed object-specific information. We observe high variance in the Summarization (Sum) accuracy due to the small number of questions in this category and the randomness of LLM. Similar improvements have been observed on the MovieChat-1K in Table 4.3, where the accuracy improved by 10.45% with action and object tokens. Ultimately, using all three token types together achieves the highest performance, demonstrating the complementary strengths of scene, action, and object tokens as a unified representation in long video understanding.
@@ 4.3 Ablation Study
Table 4: The effects of different types of tokens on LVbench. Using any type of tokens outperforms the random baseline, while applying all three types of tokens brings the best performance.
@@
We observe that the latency bottleneck mostly comes from the tracker when processing multiple objects. Table 4.3 presents an analysis of different trackers for extracting action tokens. While YOLO+BoT-SORT operates in a much higher frame rate (14 FPS), it lacks open-set detection capabilities. On the other hand, SAM2 can detect all objects within a scene but operates significantly slower (8FPS). Our findings indicate that the proposed action tokens does not require capturing all the objects in the scene. In practice, YOLO-X pretrained on COCO, with BoT-SORT, effectively captures essential information for action token extraction, maintaining a balance between performance and efficiency.
Attention learning. Table 4.3 highlights the effectiveness of query relevance and token diversity terms in Section 3.2. When $\alpha\!=\!0$, the optimization focuses on the diversity term to capture extensive contextual information for long-video tasks such as Summarization (Sum) and Reasoning (Rea), achieving reasonable results comparable to other state-of-the-art approaches in Table 1. With $\alpha\!=\!1$, the optimization prioritizes the relevance term to reduce redundancy by sampling tokens that closely align with the query, so performance can be improved for tasks like Key Information Retrieval (KIR). However, this approach loses global context and results in weaker summarization performance. Balancing both terms yields the best overall results.
@@
Table 6: Ablation studies on LVBench. SEAL defaults to global mode with YOLO for efficiency, while SAM2 performs similarly. Global mode excels in long-context tasks, and streaming mode in temporally-intensive tasks. Combining Relevance and Diversity in attention learning achieves the best performance.
Figure 4: Accuracy vs. efficiency trade-off on LVBench. SEAL runs 2-3x faster than InternVL2 at the same accuracy, and is more accurate when compared at the same FPS.
Computational Complexity. We provide an accuracy-efficiency trade-off comparison in Fig 4 where both SEAL and InternVL2 are with a varying number of tokens. FPS is calculated during inference that includes all the processing time from the raw video frames. SEAL reduces the reliance on densely sampled tokens and uses subject-level and motion aware representations/tokens to improve efficiency while maintain good accuracy. From Fig 4, at 10 FPS, SEAL is about 5% more accurate than InternVL2. When comparing at the same accuracy of 41.5% and 40%, SEAL is about 3x and 2x faster than InternVL2, respectively.
5 Conclusion and Future Work
We propose SEAL, a novel unified representation for long video understanding that addresses computational complexity, temporal redundancy, and cross-task generalization. SEAL leverages semantic decomposition to break videos into scene, object, and action tokens, reducing redundancy and enabling efficient processing. It incorporates attention learning to balance query relevance and token diversity, enhancing performance across diverse tasks. SEAL achieves state-of-the-art results on benchmarks like MovieChat-1K, LVBench, and Ego4D-NLQ, demonstrating its versatility and effectiveness for long video understanding.
Limitation. The Attention Learning module is bounded by the memory constraint for the QP solver, making it not fully end-to-end trainable. For future work, we plan to integrate Attention Learning in our streaming mode for full end-to-end learning. Developing special prediction heads to solve causal reasoning [17], where MLLM heads showed limitations, is also an interesting area for future exploration.
Acknowledgements. This work was partially supported by the Office of Naval Research (award #N00014-23-1-2417). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of ONR.
Supplementary Material
Appendix A Additional Ablations
@@
Ablation on Different $\alpha$. Figure 5 shows the performance trends across various categories for different values of $\alpha$. $\alpha=0.9$ achieves the best overall trade-off, reaching peak with the highest overall accuracy of 45.9. Conversely, extreme values like $\alpha=0.0$ or $1.0$ lead to declines in several metrics, highlighting that both diversity and relevance are essential. Therefore, $\alpha=0.9$ is the optimal choice for experiments, delivering peak performance and a well-rounded balance across all categories.
Effectiveness of Encoder for Relevance. Table A presents the relevance results computed using the BLIP (Base), CLIP (ViT-L/14) and the BLIP2 (Large) models. The results demonstrate that stronger models achieve higher effectiveness in computing relevance scores, leading to significant performance gains for SEAL.
{tabu}
to 0.47@lX[2cm c]*6X[c]@ Model Overall KIR EU Sum ER Rea TG
SEAL w/ BLIP2 45.9 51.5 41.3 39.7 47.9 43.3 32.3
SEAL w/ CLIP 42.9 48.5 38.6 36.2 46.2 36.3 32.7
SEAL w/ BLIP 40.5 41.9 38.6 39.7 40.5 47.2 32.7
Table 10: Comparison of BLIP2 with other methods on LVBench.
Appendix B Additional Results and Discussions
Comparison with LVU methods on LVBench. We provide additional comparison with LVU methods on LVBench in Table B. For a fair comparison, we follow those methods to use a 7B LLM. SEAL maintains superior performance with a much smaller LLM (7B), demonstrating the effectiveness of our proposed method.
{tabu}
to 0.47@lX[2cm c]*6X[c]@ Model Overall KIR EU Sum ER Rea TG
MovieChat [34] 22.5 25.9 23.1 17.2 21.3 24.0 22.3
TimeChat [33] 22.3 25.9 21.7 24.1 21.9 25.0 22.7
MA-LLM [10] 24.5 25.4 25.8 22.4 22.3 26.9 21.8
SEAL (7B) 36.6 44.3 33.7 27.6 36.9 32.8 30.9
Table 11: Comparison with other long video representations on LVBench.
Number of different tokens. The subset of tokens is learned as an optimization problem in Section 3.2, and the composition of tokens varies on different inputs and tasks. The averaged percentages of scene, object, and action tokens are 62.5%, 26.1%, 11.4% on LVBench, 54.3%, 25.6%, 20.1% on Moviechat, 88.5% scene tokens and 11.5% action tokens on Ego4d-NLQ. Since Ego4D-NLQ is a temporal localization task, we only utilize scene and action tokens.
Result Analysis. We evaluated the distribution of answers generated by different models, following [41], as shown in Figure 6. The Ground-Truth exhibits a fairly balanced distribution among A, B, C, and D, indicating a well-distributed dataset where no single category is disproportionately represented. MovieChat and LWM shows a dominance of category A, with significantly smaller contributions from other categories, suggesting a lack of diversity in predictions. In Gemini 1.5 Pro, the “Others” category is significantly high, indicating that Gemini 1.5 Pro produces a notable number of unrelated outputs. Our method demonstrates a distribution close to Ground-Truth, showing strong generalization and robustness.